Tech Staff Meeting Notes October 9, 2006

Current Top Priorities are:LNXT Node BB upgrade, Station Maintenance, and ET Stations. (Also wish to get INV populated.)


There are four levels of priority.
PrioritiesWish to try to get these done (or make progress on) before the next meeting (yeah right)
Second TierFit this in too if an opportunity arises. May be good to do instead of priorities but check with Mitch. Likely to be a priority in a previous and/or subsequent month
As Time PermitsFit these in when you can but not instead of first or second priorities
OtherStuff we hope to get to and don't want to forget about


This task list as of September 12, 2006. Note that everyone is still needed to work on essentially all tasks. There is however, a single person who is primarily responsible for each one. Mitch of course, is ultimately responsible for everything. Mitch is going to update this shortly.

Priorities

Lenox Agreement. Jim noticed the 12 year agreement for the LNXT node is about to expire. Mitch will renew.

BB network upgrade.
All sensors with new top-caps have the Steiner blessing. See Mitch's PDF webpage for results.
Only LNXT stations remain. LNXT cr is done. LNXT station is done. PEBM and GLAT on deck. HALT needs a 50' tower. Prefer to spend the cash and do it right rather than put up a flimsy stick.

GPS
James will coordinate with Bill or Chris when he will be going to the field and needs assistance. Notes from James:

Firstly, the radio tests in and around Tiptonville are complete. We'll likely do a relay shot from RLAP to New Markham, and then bring both sites back to the highschool in Tiptonville. We have permission from the school to use their 50' tower on-site. We'll need a second shot from our site at the Corrections Complex (NWCC) to the highschool. For this, we'll need either a roof or building mount antenna at at least 30' height. The MIS building itself (where out stuff is) can give us almost twenty feet. We'll need to buy high-gain antennas and mounting hardware, and some _serious_ lightning protection.

The second set of radio tests will be at Steele, Missouri, hopefully on 10/20. I visited this last Friday and spoke again with superintendent Mitchell Fisher. WREG has a short mast on one side of the administration building of the Steele schools for their weather station. I'm going to email our contact there to see if we can mount our antenna on that mast. The building's mechanical room has roof access and plenty of space for our radio equipment enclosure. Bob seems to be waiting for the results of this test before he buys radios and antennas.

We've not been able to secure permission to get back inside the terminal building at RLAP. To winterize the station, we need to replace the five aircells with another solar panel, battery, and a charger that will take two panels. Sure, we could trench AC power out to the station, but that would be wasted effort if we _do_ get permission to move back inside. Meanwhile, having the charger, battery, and panel hardware on-hand would be nice for one or both of our new sites around Tiptonville. Mitch and I are in agreement that we can't keep sponging off the seismic networks. We'll have to buy our own stuff.

ETN

The map shows stations with yellow triangles being new ones.
SMNC may be getting interference and the gain channel is showing up on the vertical (but doesn't look right). John will visit this week, swap the radios, and inject a known signal on the pins.
stationman-daysstatus
MGNC8done
CCNC8new
WSSR8new
TVNC8new
RBNC4done
TRYN4done
GFM4done
SMNC4done (see note above)
PLVA2abandon
BHT8move (eventually)
ETT has some breaks in the coax. Should probably replace entire length.
Flat gap, Van hill, and LRVA are on deck. Prep week of 10/23 and Chris and Greg will meet John the week of 10/30.
Need vehicle registration and insurance documents. Either the originals or something that says they are on file at State. Not a big deal in TN but sure is in other states. John as this odd aversion to jail cells. Done. Chris is looking into an insurance letter.
TZTN needs a gate and have episensors swapped. The one there now belongs to SLU.
GFM has a cable problem.
CPCT timing is repaired.

ANSS.
Bill is getting familiarized with these so he can take them over.
John upgraded gps firmware at 2 of 3 gps. Mitch doesn't recall which ones. John says Knoxville and 2 in charleston still need upgraded. This is the list from before:
SWET-3H52, ADSC-5657, and CSTN-5637 need to have the leap second corrected. SHTN-5648 and TRSC-5662 might also but need a site visit to verify.
Mitch needs to contact UKY tech to get a site visit to MVKY.
ATTN is having firewall problems. This is fixed but by accident. Probably will come round again. CVTN and ASAR optical isolation. Need to upgrade to year 3 hardware. Hardware ready. James brought back some nice pictures from CVTN to help figure out how to do it (e.g. the conduit is way too small).
SFTN and RDST need a sensor swap.

Routine Maintenance.

Unscheduled Maintenance
Jim and Bill fixed a bunch of these. Mitch isn't sure which. Some of work fine in the shop; Greg will have a look. Here's Holly's list from the last meeting:
It isn't clear what's going on with a bunch of these. They work on the bench and there aren't any measurable problems in the field. Greg will take a look.
LNXT Node:
BRGM.EHZ
PLAR.EHE & .EHN
NMAD Node:
BACM.EHE
CHRM
COKM.EHE
FLPT
POBM.EHZ
TIPT.EHZ
TOPM.EHZ
MCAM.EHE
MKTA Node:
BLAR
BOAR


Second Tier Priorities

Need to follow up on ATV training. DONE.

NCCU Gordana Vlahovic is getting a cmg6td to install at North Carolina Central University. Invoice paid. Still owe here a vault (John has it) and installation advice. Need to get a public display. Scream to Scream from the datalogger pc to the display pc would be best. Don't forget to NOT allow remote config on display scream. John will be doing this later this month. Done. Data are flowing into scream at NCCU. Al will get with Gordana to get the NCCU scream transmitting to the mudworm scream.

Steiner Brain Dump Seminar Gregs list is starting to take shape. While anyone is invited its geared to tech staff. Schedule will vary and be weather dependent to avoid field work interruptions. Will start within a few weeks.
How do we record this for posterity? Video would be enormous in size (potential of 104 hours or more of video). Audio might be useful but would not make a very good reference. Mitch thinks written is best preferably in html or pdf. Perhaps have someone (who) take everyones notes and combine them with Greg's notes and make a web page. Audio/video cache to help clarify might be useful for whoever does this. Mitch will ask the folks at the ALC what they recommend.

FDL. Mitch needs to talk with Horton about what to do about FDL laptops. Clearly we're not getting help from Bodin.
Bill is now the designated battery czar.
Bill is in charge of everything in the fdl. He'll be developing SOP's to keep track of who has what where.

3904 Power Need to draft an updated wiring diagram of 3904. Need to update the diagrams on the circuit panels. Greg put up signs on the panels with a call down list.
Greg will put an SOP for dummies on the transfer switch.


As Time Permits

Mitch needs to write an ew module to monitor the environmental channels newly installed at the nodes. Done. Its a shell script that runs twice a day. Computes several minute averages. Emails when there's a problem. Its is /gaia/home/mwithers/Projects/ENVMonitoring.

Need to come up with a good way for John to view real data (not helicorders) on is pc over dsl. Mitch will send John info on ssh and exceed.

Generators Old generator needs an exhaust pipe.

Inventory Sid has Ketchum's INV program installed on smeagol. Jim got initial training. Need to show Bill how to use it. Jim will be the lead and Bill will do initial update. Routine updates can be done by whoever does the work (requires usernames and passwords in mysql). Not sure yet how best to handle dataless seed production. Mitch needs to learn the admin tasks. Startup script is /gaia/smeagol/local/bin/runINV. Mitch will find out from Meremonte how they do site contacts, directions to station, etc, etc.


Other

Many of the air cells in storage are ancient and bad. Bill will go through them and dispose of the old. Greg has method for safe and legal disposal.

I-40 Bridge.
Boreholes are drilled and more or less secured. Need to do prep for new instrumention at HDBT and take care of drainage now that a big pile of stone and dirt was piled around the station. Greg's theory is lightning is clobbering gps. Has suppression constructed and ready to install.
Need a better pc at AZTN now that we have equivalent of two nodes plus going through it. Guralp hardware is on order. Steve and Greg will knock heads to decide on telemetry to aztn (18+ channels will exceed 950Mhz bandwidth).

NHIN.
Got the modem and internet works. Need to integrate into CERI central processing.

MPHB. Withers needs to determine current orientation so we can pull it and orient correctly. Helicorders look strange. Need to verify the data.

VBMS. Vault is done. Contractor work is complete. Alena says sensors might be available in a month but NEIC staff is currently tied up with carribean net (tough job but someone has to do it).

Langston Zack. Greg's list said to test accelerometers and mass produce hardware. This is wayback burner. Mitch will check with Chuck to verify that its even going to be done anymore. Mitch asked and its unknown. Strike from list.